# CH2I2_UV_IR_CEI_SM

### v1

@@ III Results and discussions
Figure 5: Coulomb explosion simulations of $\mathrm{CH_{2}I_{2}}$ for the $\mathrm{CH_{2}^{+}+I^{2+}+I^{2+}}$ breakup channel for different molecular geometries: the ground-state equilibrium geometry, the iso-$\mathrm{CH_{2}I_{2}}$ geometry according to Borin et al. [29], and for a specific geometry during the C-I dissociation process when the $\mathrm{CH_{2}}$-I-I angle in the dissociating molecule is the same as in iso-$\mathrm{CH_{2}I_{2}}$ due to the rotation of the $\mathrm{CH_{2}I}$ radical. The two panels show the distribution of these events (a) as a function of the sum KE of the two iodine fragments and angle between the two $\mathrm{I^{2+}}$ momentum vectors, and (b) normalized yield of the KE sum of the two $\mathrm{I^{2+}}$ ions integrated for all angles.
@@
However, from the experimental data, it is evident that the three dissociation pathways - (i) $\mathrm{CH_{2}I}$-I two-body dissociation with $\mathrm{CH_{2}I}$ rotation, (ii) $\mathrm{CH_{2}}$-I-I three-body dissociation and (iii) molecular $\mathrm{I_{2}}$ formation - could also contribute to the region with the ($\mathrm{I^{2+}}$, $\mathrm{I^{2+}}$) angle larger than 160° (at least at large delays), as seen in Fig. 4a. Contributions from these pathways need to be filtered out to identify possible signatures of isomerization to $\mathrm{CH_{2}I}$-I. The simulations in Fig. 5 show that after Coulomb explosion, the sum of the KEs of the two $\mathrm{I^{2+}}$ fragments is larger for the iso-$\mathrm{CH_{2}I_{2}}$ molecules than that of the $\mathrm{CH_{2}I_{2}}$ molecules that undergo direct dissociation into $\mathrm{CH_{2}I}$ and I via C-I cleavage. Furthermore, in the cases of three-body dissociation or molecular iodine formation, the KE of the methyl ion decreases rapidly with pump-probe delay as the $\mathrm{CH_{2}}$ fragment quickly moves away from the two iodine atoms or the $\mathrm{I_{2}}$ molecule. Therefore, the sum of the KEs of the two $\mathrm{I^{2+}}$ fragments as well as the KE of the $\mathrm{CH_{2}}$ fragment can serve as additional parameters to discriminate the different channels.
@@
In Fig. 6, the KE distribution of the methyl cation ($\mathrm{CH_{2}^{+}}$) is shown as a function of the pump-probe delay up to 800 fs and only for those events where the ($\mathrm{I^{2+}}$, $\mathrm{I^{2+}}$) angle is larger than 160°. Based on the reasoning outlined in the previous paragraph, we can conclude that the events with quickly decreasing KE, below the blue line in Fig. 6, correspond to either $\mathrm{CH_{2}}$-I-I three-body dissociation or molecular iodine formation. The events above the blue line at 13 eV mainly correspond to dissociation of $\mathrm{CH_{2}I_{2}}$ into $\mathrm{CH_{2}I}$ and $\mathrm{I}$ (with subsequent rotation of the $\mathrm{CH_{2}I}$ fragment) as well as any possible formation of iso-$\mathrm{CH_{2}I_{2}}$, with only spurious contributions from the three-body dissociation and molecular iodine formation pathways at small delays below 200 fs.
@@
Next, we select only those events with kinetic energy of $\mathrm{CH_{2}^{+}}$ ionic fragment above 13 eV from Fig. 6 and plot the sum of the KEs of the two iodine fragments as a function of pump-probe delay in Fig. 7(a). The plot reveals two contributions: an intense feature with a sum KE between 10 and 17 eV (marked by the red rectangle), and a much weaker feature at a higher sum KE, marked by the green rectangle, which contains approximately 7% of the events in the intense feature. The projection in panel (b) reveals that these two contributions (blue line) are shifted toward lower and higher energies, respectively, compared to the sum KE of the unpumped molecules (red dotted line). The integrated yield inside the green region of interest as a function of the pump-probe delay is shown as green line in Fig. 7(c). Based on the simulations shown in Fig. 5(b), we assign the events with the higher sum KE to molecular geometries resembling the iso-$\mathrm{CH_{2}I_{2}}$ structure predicted by Borin et al., which have a slightly smaller I-I distance than geometries resulting from the dissociation of $\mathrm{CH_{2}I_{2}}$. Fig. 7(c) shows that the occurrence of these structures peaks at a delay of approximately 120 fs and quickly decays again on a similar time scale. Analysis of the pump-probe data recorded at a pump wavelength of 330 nm, following the exact same steps as described above (see section II in SM), yields the blue dashed line in Fig. 7(c), which shows the same behavior as the data at 290 nm except for possibly a subtle shift towards slightly longer delays, which is at the borderline of statistical significance for the present data set.
@@
Figure 7: (a) Coincidence ion yield as a function of pump-probe delay and KE sum of the two $\mathrm{I^{2+}}$ fragments, for only those events where the angle between the two $\mathrm{I^{2+}}$ momentum vectors is greater than 160° and the KE of the $\mathrm{CH_{2}^{+}}$ fragment is greater than 13 eV (above the blue line in Fig. 6). The two regions of interest are described in the text. (b) Projection of the coincident ion yield in panel (a) on the KE axis. The corresponding KE sum distribution for "unpumped" molecules (NIR probe pulse arrives before the UV pump pulse) is shown as a red dotted line. (c) Projection of the coincident ion yield in the green ROI in (a) on the pump-probe delay axis (green), compared to the similarly analyzed yield at a pump wavelength of 330 nm (blue).
@@
Figure 8: Energy sharing between the two $\mathrm{I^{2+}}$ fragments, integrated over the pump-probe delays shown in Fig. 7, for only those events where the angle between the two $\mathrm{I^{2+}}$ momentum vectors is greater than 160° and (a) the KE of the $\mathrm{CH_{2}^{+}}$ fragment is greater than 13 eV (above the blue line in Fig. 6); or (b) the KE of the $\mathrm{CH_{2}^{+}}$ fragment is less than 13 eV (below the blue line in Fig. 6). The red and green rectangles in panel (a) correspond to events originating from the respective rectangular regions in Fig. 7.
@@
Additional information on the geometry of the transient species can be obtained by considering the energy sharing between the two detected $\mathrm{I^{2+}}$ fragments, as shown in Fig. 8. Symmetric geometries, such as the original $\mathrm{CH_{2}I_{2}}$ geometry, or any molecules excited to symmetric vibrational modes, would lead to equal energy sharing, whereas strongly asymmetric geometries such as the predicted iso-$\mathrm{CH_{2}I_{2}}$ species, would lead to a pronounced asymmetry in the energy sharing. From Fig. 8(a), it is evident that all the events surviving both the filters - large $\mathrm{I^{2+}-I^{2+}}$ angle and higher $\mathrm{CH_{2}^{+}}$ energy, have asymmetric energy sharing between the $\mathrm{I^{2+}}$ fragments, excluding, for example, the possibility that any of these events stem from a symmetric $\mathrm{I_{2}}$ formation channel. In contrast, symmetric geometries contribute strongly to the coincidence yield with low $\mathrm{CH_{2}^{+}}$ energy, as shown in Fig. 8(b), corroborating our previous conclusions that many of those events are due to $\mathrm{I_{2}}$ formation.
@@
At this point, we would like to emphasize that although detailed analysis of the CEI data clearly shows the transient formation of $\mathrm{CH_{2}I-I}$ isomer-like geometries that are consistent with previous literature [29], the CEI experiment alone is not able to confirm this as an "isomerization" process, nor can the CEI experiment identify the electronic state of the molecule when it assumes these isomer-like geometries.
@@
Whether or not the process observed and reported here can or should be referred to as "isomerization" cannot be answered by the experimental data alone and also depends on the exact definition of "isomerization" one chooses to use. Given the short time that the molecules spend in the isomer-like geometry, which is less than the timescale of the rotation of the $\mathrm{CH_{2}I}$ fragment that is produced by direct C-I bond cleavage, the process might be better described as a transient passage through isomer-like geometries rather than an actual trapping in the potential well of the isomer, but these terminologies and interpretations strongly rely on careful theoretical modeling, which is beyond the scope of the present paper.
@@ IV Conclusion
We have applied time-resolved three-body Coulomb explosion imaging to study the UV-induced (290-nm and 330-nm) photochemical dynamics in gas-phase $\mathrm{CH_{2}I_{2}}$. In addition to the primary dissociation into $\mathrm{CH_{2}I+I}$ via direct cleavage of the C-I bond, the formation of molecular $\mathrm{I_{2}}$, and the three-body dissociation into $\mathrm{CH_{2}+I+I}$ (where the latter two processes are attributed to the absorption of more than one UV photon), we also observe the transient molecular configurations resembling iso-$\mathrm{CH_{2}I_{2}}$ geometries, which are formed within approximately 100 fs after the photoexcitation and completely decay within an additional 100 fs. Within the uncertainties of the experimental data, the relative yield, formation and decay time of the transient iso-$\mathrm{CH_{2}I_{2}}$ products appear to be similar at both excitation wavelengths employed in this study, suggesting that their formation does not depend strongly on the excitation wavelength.

### v2

@@ III Results and discussions
Figure 6: Coulomb explosion simulations of $\mathrm{CH_{2}I_{2}}$ for the $\mathrm{CH_{2}^{+}+I^{2+}+I^{2+}}$ breakup channel for different molecular geometries: the ground-state equilibrium geometry, the iso-$\mathrm{CH_{2}I_{2}}$ geometry according to Borin et al. [10], and for a specific geometry during the C–I dissociation process when the $\mathrm{CH_{2}-I-I}$ angle in the dissociating molecule is the same as in iso-$\mathrm{CH_{2}I_{2}}$ due to the rotation of the $\mathrm{CH_{2}I}$ radical. The two panels show the distribution of these events (a) as a function of the KE sum of the two iodine fragments and angle between the two $\mathrm{I^{2+}}$ momentum vectors, and (b) normalized yield of the KE sum of the two $\mathrm{I^{2+}}$ ions integrated for all angles. To emphasize the structure associated with iso-$\mathrm{CH_{2}I_{2}}$ and its overlap with the other distributions, a white contour representing the region containing 90% of the iso-$\mathrm{CH_{2}I_{2}}$ event density is overlaid on panel (a).
@@
However, from the experimental data, it is evident that the three dissociation pathways—(i) $\mathrm{CH_{2}I}$-I two-body dissociation with $\mathrm{CH_{2}I}$ rotation, (ii) $\mathrm{CH_{2}}$-I-I three-body dissociation and (iii) molecular $\mathrm{I_{2}}$ formation—could also contribute to the region with the ($\mathrm{I^{2+}}$, $\mathrm{I^{2+}}$) angle larger than 160 (at least at large delays), as seen in Fig. 5(a). Contributions from these pathways need to be filtered out to identify possible signatures of isomerization to $\mathrm{CH_{2}I}$–I. The simulations in Fig. 6 show that after Coulomb explosion, the sum of the KEs of the two $\mathrm{I^{2+}}$ fragments is larger for the iso-$\mathrm{CH_{2}I_{2}}$ molecules than that of the $\mathrm{CH_{2}I_{2}}$ molecules that undergo direct dissociation into $\mathrm{CH_{2}I}$ and I via C–I cleavage. Furthermore, in the cases of three-body dissociation or molecular iodine formation, the KE of the methyl ion decreases rapidly with pump-probe delay as the $\mathrm{CH_{2}}$ fragment quickly moves away from the two iodine atoms or the $\mathrm{I_{2}}$ molecule. Therefore, the sum of the KEs of the two $\mathrm{I^{2+}}$ fragments as well as the KE of the $\mathrm{CH_{2}}$ fragment can serve as additional parameters to discriminate the different channels.
@@
In Fig. 7, the KE distribution of the methyl cation ($\mathrm{CH_{2}^{+}}$) is shown as a function of the pump-probe delay up to 800 fs and only for those events where the ($\mathrm{I^{2+}}$, $\mathrm{I^{2+}}$) angle is larger than 160. Based on the reasoning outlined in the previous paragraph, we can conclude that the events with quickly decreasing KE, below the blue line in Fig. 7, correspond to either $\mathrm{CH_{2}}$-I-I three-body dissociation or molecular iodine formation. The events above the blue line at 13 eV mainly correspond to dissociation of $\mathrm{CH_{2}I_{2}}$ into $\mathrm{CH_{2}I}$ and $\mathrm{I}$ (with subsequent rotation of the $\mathrm{CH_{2}I}$ fragment) as well as any possible formation of iso-$\mathrm{CH_{2}I_{2}}$, with only spurious contributions from the three-body dissociation and molecular iodine formation pathways at small delays below 200 fs. The choice of 13 eV as the boundary is motivated by the experimental data (Fig. S5), where a distinct separation is observed between the decaying low-KE band and the high-KE island in the 100–200 fs region. Although the constant 13 eV horizontal line intersects other features at later delays, our analysis focuses only on the early-delay (< 300 fs) region, so those intersections do not affect the conclusions.
@@
Next, we select only those events with kinetic energy of $\mathrm{CH_{2}^{+}}$ ionic fragment above 13 eV from Fig. 7 and plot the sum of the KEs of the two iodine fragments as a function of pump-probe delay in Fig. 8(a). The plot reveals two contributions: an intense feature with a KE sum between 10 and 17 eV (marked by the red rectangle), and a much weaker feature at a higher KE sum, marked by the green rectangle, which contains approximately 7% of the events in the intense feature. The projection in panel (b) reveals that these two contributions are shifted toward lower (solid line, red section) and higher energies (solid line, green section), respectively, compared to the KE sum of the unpumped molecules (black dotted line). The integrated yield inside the green region of interest as a function of the pump-probe delay is shown as green line in Fig. 8(c). Based on the simulations shown in Fig. 6(b), we assign the events with the higher KE sum to molecular geometries resembling the iso-$\mathrm{CH_{2}I_{2}}$ structure predicted by Borin et al., which have a slightly smaller I–I distance than geometries resulting from the dissociation of $\mathrm{CH_{2}I_{2}}$. To quantify the change in the I–I distance, we calibrate the relationship between the KE sum of the two I2+ ions and the I–I distance using Coulomb explosion simulations with varied I–I distances. Applying this calibration to the experimental data yields an estimated I–I distances of $\sim$3.0 Å for the events with high KE sum and $\sim$4.6 Å for the events with low KE sum (see Fig. S8), compared to $\sim$3.58 Å in the equilibrium geometry. Fig. 8(c) shows that the occurrence of these iso-$\mathrm{CH_{2}I_{2}}$-like structures peaks at a delay of approximately 120 fs and quickly decays again on a similar time scale. Analysis of the pump-probe data recorded at a pump wavelength of 330 nm, following the exact same steps as described above (see Sec. III in SM), yields the blue dashed line in Fig. 8(c), which shows the same behavior as the data at 290 nm except for possibly a subtle shift towards slightly longer delays, which is at the borderline of statistical significance for the present data set. We focus on the 290 nm results (which have superior statistics and signal-to-noise ratio due to the higher absorption cross section at 290 nm[25]) in the main text, while noting that the 330 nm data, which is shown in more detail in the SM, exhibit qualitatively similar dynamics, as can be seen in Fig. 8(c).
@@
Figure 8: (a) Coincidence ion yield as a function of pump-probe delay and KE sum of the two $\mathrm{I^{2+}}$ fragments, for only those events where the angle between the two $\mathrm{I^{2+}}$ momentum vectors is greater than 160° and the KE of the $\mathrm{CH_{2}^{+}}$ fragment is greater than 13 eV (above the blue line in Fig. 7). The two regions of interest are described in the text. (b) Projection of the coincident ion yield in panel (a) on the KE axis. The corresponding KE sum distribution for "unpumped" molecules (NIR probe pulse arrives before the UV pump pulse) is shown as a black dotted line. (c) Projection of the coincident ion yield in the green ROI in (a) on the pump-probe delay axis (green), compared to the similarly analyzed yield at a pump wavelength of 330 nm (blue).
[…]
@@
Figure 10: Energy sharing between the two $\mathrm{I^{2+}}$ fragments, integrated over the pump-probe delays shown in Fig. 8, for only those events where the angle between the two $\mathrm{I^{2+}}$ momentum vectors is greater than 160 and (a) the KE of the $\mathrm{CH_{2}^{+}}$ fragment is greater than 13 eV (above the blue line in Fig. 7); or (b) the KE of the $\mathrm{CH_{2}^{+}}$ fragment is less than 13 eV (below the blue line in Fig. 7). The red and green rectangles in panel (a) correspond to events originating from the respective rectangular regions in Fig. 8.
[…]
@@
Additional information on the geometry of the transient species can be obtained by considering the energy sharing between the two detected $\mathrm{I^{2+}}$ fragments, as shown in Fig. 10. Symmetric geometries, such as the original $\mathrm{CH_{2}I_{2}}$ geometry, or any molecules excited to symmetric vibrational modes, would lead to equal energy sharing, whereas strongly asymmetric geometries such as the predicted iso-$\mathrm{CH_{2}I_{2}}$ species, would lead to a pronounced asymmetry in the energy sharing.
[…]
@@
At this point, we would like to emphasize that although detailed analysis of the CEI data strongly suggests the transient formation of CH2I–I isomer-like geometries that are consistent with previous literature [10], the mapping from position to fragment momentum space in CEI is not unique [57], and we cannot exclude that other geometries would lead to similar CEI signatures. However, we consider such isomer-like geometries to be the most likely reason for our observations, given the geometric constraints imposed by the various observables we considered in conjunction with the observed delay-dependent evolution. That said, it also needs to be stated that the CEI experiment alone is not able to confirm this as an "isomerization" process, nor can the CEI experiment identify the electronic state of the molecule when it assumes these isomer-like geometries.
@@
Whether or not the process observed and reported here can or should be referred to as "isomerization" cannot be answered by the experimental data alone and also depends on the exact definition of "isomerization" one chooses to use. Given the short time that the molecules spend in the isomer-like geometry, which is less than the timescale of the rotation of the $\mathrm{CH_{2}I}$ fragment that is produced by direct C–I bond cleavage, the process might be better described as a transient passage through isomer-like geometries rather than an actual trapping in the potential well of the isomer, but these terminologies and interpretations strongly rely on careful theoretical modeling, which is beyond the scope of the present paper.
@@ IV Conclusion
We have applied time-resolved three-body Coulomb explosion imaging to study the UV-induced (290 nm and 330 nm) photochemical dynamics in gas-phase $\mathrm{CH_{2}I_{2}}$. In addition to the primary dissociation into CH2I + I via direct cleavage of the C–I bond, the formation of molecular $\mathrm{I_{2}}$, and the three-body dissociation into CH2 + I + I (where the latter two processes are attributed to the absorption of more than one UV photon), we also observe the transient molecular configurations resembling iso-$\mathrm{CH_{2}I_{2}}$ geometries, which are formed within approximately 100 fs after the photoexcitation and completely decay within an additional 100 fs. Within the uncertainties of the experimental data, the relative yield, formation and decay time of the transient iso-$\mathrm{CH_{2}I_{2}}$-like products appear to be similar at both excitation wavelengths employed in this study, suggesting that their formation does not depend strongly on the excitation wavelength. While the present experiment cannot accurately quantify the branching ratio of these transient geometries since the strong-field ionization rate is strongly dependent on the electronic state and molecular geometry, it appears that only a small fraction of excited molecules form isomer-like geometries. Therefore, it is conceivable that other ultrafast structural studies, such as the gas-phase UED experiment by Liu et al.[45], may not have been able to observe these structures since it lacked the required signal-to-noise ratio to discern extremely weak channels and also did not have sufficient temporal resolution to resolve such short-lived transient structures.
V Supplementary Material
Supplementary material includes probe-only spectra, supporting results (Dalitz plots, energy-sharing ratios, and I–I distance estimate), comparison results from 330 nm pump, UV power-dependence measurements, and a movie of the delay evolution.